You have 10 uM (1*10^-6M) of Tartaric Acid. pKa1=2.89, pKa2=4.4

Calculate pH
Answer #1


ОН О НО он о но Yон + нон о ён Ка1 — Y он * H50* o Он он о Ka2 « он о но Аое , нон - о он ... ~ . . . о он ..................

pKaj = 2.89 Kaj = 10-2.89 = 1.288 x 10-03 pKa2 = 4.4 Ka2 = 10-4.4 = 3.98 x 10-05 Kaj >> Kaz Kai = [A-][H3O+] Kai = [A] [A-] ~